OpenStudy (anonymous):

Solve 3x2 = 12x. x = 0, x = 4 x = 0, x = −4 x = 3, x = 4 x = 3, x = −4

OpenStudy (anonymous):

is that 3x^2=12x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3x2 = 12x

OpenStudy (anonymous):

so it's 3*x*2=12*x?

OpenStudy (anonymous):

or 3*2=12*x

OpenStudy (anonymous):

x = 0, x = 4

OpenStudy (anonymous):

Could you make the equation more explicit? (use x only for the variable x and * for multiplying)

OpenStudy (anonymous):

or draw it.

OpenStudy (anonymous):